A patient in a hospital is given soup daily ina cylinderical bowl of diameter 7cm.If the bowl is filled with soup to a height of 4cm,how much soup the hospital daily to serve 250 patients?
 
 

Grade:9

1 Answers

Arun
6 years ago
Dear Nirmal
 

Radius (r) of cylindrical bowl = 7/2 = 3.5 cm

Height (h) of bowl, up to which bowl is filled with soup = 4 cm

Volume of soup in 1 bowl = πr2h

 = 22/7 * (3/5)2 *4

= (11 × 3.5 × 4) cm3

= 154 cm3

Volume of soup given to 250 patients = (250 × 154) cm3

= 38500 cm3

= 38.5 litres.

 

Regards

Arun (AskIITians Forum Expert)